### v1

# Integrated Positioning and Communication via LEO Satellites: Opportunities and Challenges

Abstract
Low Earth orbit (LEO) satellites, as a prominent technology in the 6G non-terrestrial network, offer both positioning and communication capabilities. While these two applications have each been extensively studied and have achieved substantial progress in recent years, the potential synergistic benefits of integrating them remain an underexplored yet promising avenue. This article comprehensively analyzes the integrated positioning and communication (IPAC) systems on LEO satellites. By leveraging the distinct characteristics of LEO satellites, we examine how communication systems can enhance positioning accuracy and, conversely, how positioning information can be exploited to improve communication efficiency. In particular, we present two case studies to illustrate the potential of such integration. Finally, several key open research challenges in the LEO-based IPAC systems are discussed.
Index Terms:
LEO satellite, positioning, communication, IPAC, 5G, 6G.
I Introduction
With the rise of applications such as the internet of everything, self-driving vehicles, and spatial intelligence, the demand for accurate positioning and reliable communication continues to grow. As a pivotal technology, low Earth orbit (LEO) satellites nowadays are playing an important role in the fields of mobile radio positioning and communication in fifth generation (5G) and beyond. In this section, we begin by introducing the unique features of LEO satellite systems, followed by a brief review of their positioning and communication capabilities.
I-A LEO Features
[…]
leo satellites are positioned at orbits altitudes ranging from 160 km to 2,000 km. Compared to medium Earth orbit (MEO) and geostationary orbit (GEO) satellites, their lower orbital altitude significantly reduces the signal propagation distance, which results in lower path loss and thus stronger received signals, as well as reduced delay. Furthermore, the wide distribution and large number of LEO satellites enable global coverage and consistent visibility to ground users. Notable constellations such as Starlink, OneWeb, and Telesat, which comprise hundreds to thousands of satellites, form effective positioning and communication networks. Notably, recent technological advancements have also significantly reduced the costs associated with the satellites’ fabrication and launching. The advancement of smaller, lighter microsatellites, requiring fewer materials, has facilitated the use of more cost-effective launch vehicles. Additionally, the utilization of Reusable Launch Vehicles (RLVs) has further reduced costs and improved the efficiency of space launches. The quick turnaround ability of RLVs also allows for a substantial increase in launch frequency [1].
Typically, LEO satellites can operate in two different paradigms: the ground-based network architecture and the space-based network architecture, as detailed in 3GPP TR 23.700 [2] and illustrated in Fig. 1.
I-A1 Ground-Based Network Architecture
In the ground-based network architecture, the satellites use a transparent payload, meaning the satellite transponders are transparent to signals [3]. In this scenario, satellites serve to connect users and base stations by merely forwarding, filtering, and amplifying signals without decoding/interpreting them. They receive signals from the ground and forward them back to Earth. The user sends data through the service link, i.e., the radio link from the user equipment (UE) to the satellite. After being forwarded by satellites, the data is transmitted through feeder link from the satellite network to the terrestrial system. Finally, the data is directed to the end user or the data processing center through the terrestrial data network.
I-A2 Space-Based Network Architecture
Unlike the ground-based network where satellites simply relay signals, the space-based network architecture enables more complex communication tasks. The space segment often utilizes regenerative payloads, which not only perform the functions of transparent payloads but also understand and process signals, such as coding/decoding and modulation/demodulation [3]. Additionally, there exist inter-satellite links connecting these regenerative payloads, thereby enabling collaborative designs among the satellites with regard to both positioning and communication.
Figure 1: Structure of LEO satellite systems, divided into the space, user, and ground segments. The space segment of the system provides service and feeder links with various payload types, including regenerative and transparent payloads. The user segment encompasses ground users, infrastructures, and building environments. The ground segment consists of gateways, base stations, data centers, core networks, and other related components.
I-B LEO Satellite-Based Positioning
In recent years, positioning based on LEO satellites, which forms a part of the broader framework of LEO satellite positioning, navigation, and timing (PNT), has attracted significant attention for its potential to complement global navigation satellite system (GNSS) or serve as a reliable backup in case GNSS systems fail. LEO PNT can be implemented through several strategies [4]. One effective strategy employs dedicated LEO satellites, which are specifically engineered to deliver optimal PNT services. Alternatively, LEO satellites not initially designed for PNT can still be utilized opportunistically to gather PNT data through signals of opportunity (SoOP). In such systems, LEO satellite-based communication systems can be enhanced by PNT services. This enhancement can be achieved by either integrating a PNT-specific payload into existing communication satellites or by incorporating PNT functionalities directly into the physical layer of satellite communication systems.
Positioning with LEO satellites provides global coverage and offers substantial advantages over systems that utilize satellites in higher orbits, such as MEO and GEO. The proximity of LEO satellites to Earth results in higher received signal strength, which not only enhances positioning accuracy but also facilitates effective indoor signal penetration where GNSS fails to reach. This robust signal strength also bolsters security by offering strong resistance to interference, including jamming and spoofing attacks. Additionally, the fast-moving nature of LEO satellites allows the utilization of Doppler shift for precise positioning [4]. Moreover, the significantly larger constellations of LEO satellites, outnumbering GNSS constellations by several orders of magnitude, have the potential to offer higher positioning accuracy through increased observable redundancy and expanded coverage, particularly in urban environments with open-sky conditions.
I-C LEO Satellite-Based Communication
The main differences between LEO communication and terrestrial communication lie in the severer path loss, the greater propagation delay, and the stronger Doppler effect.
TABLE I: The Large-Scale Path Loss of the Satellite-Terrestrial Wireless Channel [3, 5]
Item Features
Free space path loss Depends on the separation distance and signal frequency.
Shadow fading loss Depends on the geometric elevation angle, building intensity, and signal frequency. Usually modeled as the normal distribution.
Clutter loss Depends on the geometric elevation angle and signal frequency. In the line-of-sight (LoS) condition, this term is negligible and should be set to 0 dB.
Attenuation due to atmospheric absorption Depends on elevation angle, altitude above sea level, water vapor density (absolute humidity), and signal frequency. It can generally be disregarded at frequencies below 10 GHz. However, for elevation angles less than 10 degrees, it is advisable to account for it at any frequency above 1 GHz.
Attenuation due to either ionospheric or tropospheric scintillation Corresponds to rapid fluctuations of the received signal amplitude and phase. Scintillations depend on location, time-of-day, season, solar and geomagnetic activity. Ionosphere scintillation shall only be considered for frequencies below 6 GHz, while tropospheric propagation shall only be considered for frequencies above 6 GHz.
Building penetration loss It varies significantly depending on the location within the building and the construction details of the building. Modern, thermally-efficient buildings with materials such as metalized glass and foil-backed panels generally exhibit significantly higher penetration loss than ‘traditional’ buildings.
The path loss model of LEO satellite channel typically consists of two components: the large-scale path loss and the small-scale fading. According to 3GPP TR 38.811 [3], the large-scale path loss of satellite-terrestrial channels comprises several components: free space path loss, shadow fading loss, clutter loss, atmospheric absorption, ionospheric/tropospheric scintillation, and potential building penetration loss. A summary of these attenuation effects and their key characteristics is listed in Table I. In addition to large-scale path loss, small-scale fading, or fast fading, must be considered. Most satellite channel models rely on flat fading. The statistical model outlined in [6] for mixed propagation conditions in the land mobile-satellite service is one of the most up-to-date references. Detailed modeling methodology of frequency-selective fading models is described in Section 7.5 of 3GPP TR38.901[7] and Section 6.7.2 of 3GPP TR38.811[3]. These models indicate that the path loss in satellite-terrestrial links is significantly higher than in terrestrial networks due to not only longer propagation distances but also a complex atmospheric environment.
The long propagation distance of satellite channels also results in substantial propagation delays, which increases communication latency. Additionally, compared to GNSS, the lower orbital altitude of LEO satellites causes more significant variations in propagation distance to different ground users, leading to highly diverse propagation delays from different users. Furthermore, considering the fast movement of LEO satellites, these propagation delays are also time-varying. This effect can potentially hinder the usability of delay-sensitive services. Consequently, some delay-related metrics like delay satisfaction ratio and delay variance are adopted as key performance indicators of the overall LEO satellite network.
Another unique feature of LEO satellite communication is the strong Doppler frequency shift due to the high-speed mobility of LEO satellites. Typically, the maximum Doppler shift in satellite signals increases with higher operation frequencies, faster satellite velocities stemming from lower orbital altitudes, and smaller elevation angles [8]. For instance, a LEO satellite operating at 30 GHz with an orbital altitude of 800 km exhibits a maximum Doppler frequency shift of approximately 660 kHz. While the Doppler effect can be beneficial for positioning, it is generally harmful for communication, and a Doppler compensation step is usually required to maintain signal integrity and communication quality. Furthermore, according to basic wireless communication principles, the channel coherence time is approximately the inverse of the channel’s Doppler spread. This implies that the coherence time of LEO satellite channels will be much shorter than that of terrestrial channels due to the stronger Doppler shift.
II Why LEO Positioning and Communication Should Be Integrated?
While LEO satellites possess promising positioning and communication capabilities, these two functions are often discussed separately. We argue that considerable mutual enhancement can be achieved through their synergy. In the following subsections, we will demonstrate the potential gains from such integrated positioning and communication (IPAC) system from both communication and positioning viewpoints.
II-A The Role of Positioning in Enhancing Communication
In this subsection, we will demonstrate the critical role of position information in enhancing LEO communication performance through several specific applications, based on the unique features of LEO communication channels outlined in Section I-C.
II-A1 Position Information-Enabled Fast Beamforming
With the ability to concentrate transmit power in targeted directions, multi-antenna beamforming is crucial in combating severe path loss and enhancing LEO satellite communication capacity, just as it does in terrestrial networks. However, beamforming techniques typically require channel state information (CSI). As analyzed in the previous subsection, LEO satellite channels are facing the unique challenge of short coherence time and fast variation. This imposes significant challenges for real-time channel estimation and beamforming. Considering the LoS-dominant propagation and the known satellite orbits, CSI-free beamforming that leverages positioning information represents a potential solution. Given the locations of the users and the motion of the satellites, directional beamformers can be designed predictively and dynamically towards the LoS direction. This approach holds great promise for avoiding the high time cost associated with channel estimation. However, beamforming based on the LoS channel only may result in a lack of diversity, which can undermine the reliability of LEO satellite communication systems. With the increasingly dense deployment of LEO satellites, multi-satellite-enabled cooperative beamforming, which potentially enriches diversity and boosts received power, can be a promising solution.
II-A2 Position Information-Based Fast Timing Advance Update
In 5G new radio (NR) systems, one process that is significantly impacted by the large propagation delay is the random access procedure. To reduce timing misalignment interference, 5G NR employs a timing advance scheme for random access. Specifically, a designed time offset is applied to each uplink signal to make sure that the BS can receive the uplink frames from different users within the range of a single cyclic prefix (CP) [9]. Considering the rapid drift and high variance of the propagation delays in LEO channels, very fast update of the timing advance is necessary to reach this standard. To address this issue, 3GPP TR 38.811 [3] has already highlighted that leveraging knowledge of UE locations and satellite ephemeris can be a promising solution. Given the position information of UEs, the delays can be easily computed and predicted, thus enabling fast timing advance update. Initial studies have demonstrated that by effectively utilizing UE position information, a 5G NR-adapted timing advance estimation scheme can be implemented [10].
II-A3 Other Benefits
Apart from the benefits to beamforming and timing advance update, position information can also improve other aspects of communication. For example, acquiring accurate UE position can be critical for Doppler compensation of LEO satellite signals, which has not yet been studied much. In addition, the accurate position information of both satellites and UEs is crucial for routing optimization (or path planning) in indirect satellite communications through multi-hop connectivity, which is a key challenge in massive LEO constellations. Overall, position information holds significant potential for enhancing LEO satellite communication performance. Arguably, it plays a more pronounced role in LEO communication systems than in terrestrial networks due to the distinctive characteristics of LEO satellite systems.
II-B The Role of Communication in Enhancing Positioning
Not only does positioning benefit communication, but communication also enhances positioning capabilities. This subsection explores various perspectives on how communication, particularly in the context of LEO satellite systems, aids in positioning.
II-B1 Infrastructure Perspective
Owing to using higher frequency bands, LEO satellite communication systems can employ advanced infrastructures such as phased array antennas and hybrid beamforming structures to enhance communication performance, which also benefit positioning applications significantly. For instance, these infrastructures enable the acquisition of the angle-of-departure (AoD) and angle-of-arrival (AoA) observations, which are unavailable in GNSS that uses single-antenna transmitters and receivers. These angle observations are closely related to the UE position and can thus be explored to boost positioning performance. Moreover, when massive multiple-input multiple-output (MIMO) technology is deployed on LEO satellites, the accuracy of these angle observations can be further improved because of the higher spatial resolution. Besides, in 5G/ sixth generation (6G) systems, the extensive terrestrial communication infrastructure can be leveraged and integrated with satellite-based positioning systems to enhance localization capabilities. Traditional GNSS applications already demonstrate the use of ground stations to assist high-precision positioning, as seen in real-time kinematic (RTK) positioning systems [11]. In emerging LEO satellite systems, theoretical studies have also shown that integrating ground-based technologies, such as reconfigurable intelligent surface (RIS), have the potential to significantly enhance positioning and tracking accuracy, particularly in complex suburban and urban environments [12].
II-B2 Network Coordination Perspective
In addition to the advantages provided by new communication infrastructures, efficient inter-satellite and satellite-to-ground communication can also significantly enhance positioning performance by enabling cooperation across the network. For example, multiple LEO satellites equipped with MIMO antenna arrays can achieve superior performance through coordinated beamforming, as opposed to independent beamforming schemes. However, this coordination is only feasible with efficient inter-satellite communication links. Furthermore, it is preferable for the signals from each satellite to be orthogonal (in time or frequency), to mitigate mutual interference and ensure accurate positioning, which requires communication and coordination between satellites. Therefore, information sharing between different positioning anchors (satellites or BSs), enabled by efficient communication links, is the key to enhancing the positioning performance of the entire system.
II-B3 Other Benefits
LEO communication technologies enable satellites to synchronize their clocks with ground-based atomic clocks via signals, ensuring highly accurate time data for position calculations. Additionally, LEO communication technologies facilitate the implementation of fault-tolerant systems capable of automatically switching control and rerouting data in the event of satellite or ground station failures. This redundancy is essential for maintaining uninterrupted positioning services. Furthermore, the two-way communication capability of LEO systems enhances dynamic and responsive positioning services, providing an improvement over systems lacking a return channel.
III Case Studies
To intuitively demonstrate the superiority of IPAC in LEO satellite systems, this section presents two detailed case studies. Through numerical simulations, we first illustrate how UE position information can enhance LEO communication performance, and then, conversely, show how communication links can improve positioning performance.
III-A Simulation Setup
Considering a downlink scenario, we simulate LEO satellites at an altitude of 400 km, each with a transmit power of 60 dBm and a 240 MHz bandwidth, operating at 28 GHz. The ground UE is modeled as a single-antenna receiver with a noise power spectral density (PSD) of -174 dBm/Hz. The wireless channels are generated using the Rician fading model, with a Rician factor of 3 by assuming the user is in a rural environment [12, Table II]. By default, the LEO satellites are equipped with a $20\times 20$ antenna array with half-wavelength spacing, oriented towards the Earth. For simplicity, we employ an analog beamforming schemes.
III-B Spectral Efficiency Evaluation
Figure 2: Comparison of the two beamforming methods for spectral efficiency. Note that the channel estimation error here is defined as the ratio between the error variance and the squared channel gain.
Fig. 2 evaluates the spectral efficiency of the communication between a single LEO satellite and a ground UE. We test two different beamforming schemes: (i) Beamforming using an outdated channel estimate with an estimation error, and (ii) beamforming using a reconstructed LoS channel based on the UE location with an uncertainty. As mentioned earlier, the rapid movement of LEO satellites significantly reduces the coherent time of LEO channel. Consequently, traditional channel estimation methods fall short, as the estimated channel can quickly become outdated. Existing studies show that the none-line-of-sight (NLoS) component is more significantly affected [13]. Based on this, scheme (i) models the outdated channel as a channel having the same LoS component but with a randomly different NLoS component compared to the true channel at the transmission moment. In addition, we assume the obtained outdated channel contains a Gaussian estimation error. In contrast, scheme (ii) reconstructs the LoS component of the channel based on the UE position and the satellite ephemeris, and then forms the directional beam accordingly. This method circumvents the time-consuming channel estimation step. Similarly, the prior UE location information is assumed to have Gaussian uncertainty.
As shown in Fig. 2, the performance of both methods improves as the channel estimation error or UE location uncertainty decreases. However, the superiority of the UE location-based beamforming is clearly evident. As highlighted by a red line, the UE location-based beamforming with 10 km location uncertainty achieves a spectral efficiency comparable to that of the outdated channel-based beamforming with a 0.1 % channel estimation error. Clearly, achieving a 10 km positioning accuracy is much easier than reducing channel estimation error to within 0.1 %. Under the considered setup, when positioning accuracy is better than 10 km, which is typically realizable, the UE location-based beamforming consistently outperforms the outdated channel-based beamforming. This indicates that, given the issue of channel aging, leveraging UE location is critical for providing more practical and effective beamforming solutions and enhancing communication spectral efficiency.
III-C Positioning Error Bound Evaluation
Next, we illustrate how communication can enhance LEO positioning accuracy by employing the Cramér-Rao Bound (CRB) as a performance metric. As previously discussed, the enhancement of localization through communication involves two key aspects: advanced infrastructure and network-level collaboration. For infrastructure, we compare the positioning CRB between scenarios using satellites equipped with a single antenna (typical of traditional GNSS setups) and those with an antenna array (common in high-frequency communication systems). The antenna array enables the acquisition of AoD observations, which are not available with single-antenna transmitters. For network-level collaboration, we evaluate cooperative versus non-cooperative positioning scenarios. Through inter-satellite communication, cooperative satellites can transmit orthogonal signals to the user, thus minimizing interference. In contrast, non-cooperative satellites transmit signals independently, leading to stronger mutual interference.
[…]
Figure 3: Positioning CRB under different system setups: (i) Single-antenna cooperative (SA-C) satellites. (ii) Multi-antenna cooperative (MA-C) satellites. (iii) Multi-antenna non-cooperative (MA-NC) satellites. In cooperative scenarios, signals from different satellites are orthogonal, enabling independent separation in the receiver. In non-cooperative cases, however, these signals are superposed into a single set of observations. Besides, each satellite deploys an $N\times N$ square antenna array, with $S$ representing the number of satellites.
Fig. 3 presents the calculated positioning CRB across antenna arrays ranging from $2\times 2$ to $32\times 32$ elements. In the single-antenna scenarios, the positioning error bound remains consistently high. However, when antenna arrays are employed (thus the AoD observation is available), the CRB is significantly reduced and decreases as the number of antennas increases. Moreover, it is observed that cooperative positioning with orthogonal signal transmission from each satellite substantially outperforms the non-cooperative cases. For example, in the cooperative positioning scenarios, 4 satellites achieve a lower positioning bound than 5 satellites in the non-cooperative transmission case. This indicates that orthogonal signal transmission enabled by inter-satellite communication can optimize resource utilization and achieve better positioning accuracy even with fewer satellites. These findings reveal that the combination of advanced antenna hardware and effective inter-satellite communication holds the potential to improve the overall performance of LEO satellite positioning systems.
IV Open Problems
Despite its promising potential, the IPAC system on LEO satellites still faces numerous unresolved challenges. This section summarizes several key open problems in the field.
IV-A Doppler Shift
The strong Doppler effect on LEO satellite systems is a double-edged sword. On the one hand, the Doppler shift provides valuable information about the relative velocity between users and satellites, which is beneficial to positioning. On the other hand, this effect can be harmful to communication. As mentioned in Section I-C, an approximate 700 kHz Doppler frequency is comparable to the communication symbol rate. This will lead to severe inter-subcarrier interference and further degrade the demodulation performance. Additionally, the high Doppler shift variation in LEO systems may exceed the tolerances that a communication system can support [14]. Thus, advanced algorithms are required to balance the utilization of the Doppler effect for accurate positioning and the mitigation of its negative effects in communication.
IV-B TDD vs. FDD
The LEO IPAC system can operate in both frequency division duplexing (FDD) and time division duplexing (TDD) modes, as illustrated in Fig. 4. Note that here duplexing extends beyond just the uplink and downlink transmissions; it also applies to positioning and communication functions. In an integrated system with a complex network structure, varying user requirements, and diverse component characteristics, efficiently allocating spectrum and time resources between positioning pilot symbols and communication data streams is critical. Typically, TDD is advantageous for systems requiring dynamic adaptability and efficient spectrum use but falls short in systems requiring extreme real-time performance. In contrast, FDD excels in scenarios demanding real-time IPAC collaboration and stable performance but at the cost of increased spectrum usage and hardware complexity. The appropriate selection and adjustment of TDD and FDD to meet diverse requirements is a significant challenge in LEO IPAC systems.
Figure 4: Frequency division vs. time division duplexing schemes for positioning and communication.
IV-C Satellite Orbit Errors
[…]
Figure 5: Tested positioning RMSE and theoretical error bounds versus channel observation accuracy under various levels of satellite position mismatch. Here, we consider a positioning system based on observations including channel delay, Doppler shift, and AoD. The RMSEs are evaluated using a maximum likelihood estimator.
The orbit information of LEO satellites is typically not accurately known due to several factors. Perturbing forces, such as atmospheric drag and solar radiation pressure, are primary contributors to these inaccuracies. Moreover, the parametrization of the orbit stored in two-line element (TLE) files further complicates precise orbit prediction [4]. These factors introduce orbital errors, causing mismatches in the satellites’ position and orientation information, which ultimately degrade positioning and communication performance. Fig. 5 presents an evaluation of the positioning RMSE versus channel observation accuracy (using channel delay as a representative) under different levels of satellite position mismatch. The results indicate that such mismatches can lead to performance saturation, limiting the achievable accuracy of the system. A similar negative impact can be observed in communication performance, especially in an IPAC system.
IV-D Satellites Handover
Satellite handover refers to the process of transferring the communication link from one satellite to another as they traverse the coverage area of a ground station or the signal reception range of UE. Due to their rapid movement relative to the Earth’s surface, LEO satellites require frequent handovers to ensure continuous communication and accurate positioning. This regular switching intensifies the complexity of network resource management and challenges the maintenance of uninterrupted service continuity. Managing handovers within a constellation of potentially hundreds or thousands of satellites demands advanced network management systems and sophisticated algorithms to coordinate traffic and enhance network performance.
IV-E Resource Allocation to Multi-Users
Given that thousands of satellites are simultaneously serving numerous users on Earth, resource management remains a significant challenge. Joint time-frequency-space resource allocation needs optimization, which requires substantial computational resources. When integrating positioning and communication, additional considerations in resource allocation are necessary based on users’ specific needs. For instance, positioning reference signals (PRSs) defined in 3GPP TS 38.211 [15] can be integrated into orthogonal frequency-division multiplexing (OFDM) time-frequency grids as positioning pilots. Users requiring accurate positioning services should receive more PRSs, while those prioritizing communication services should have more resources allocated to data transmissions. Moreover, resource allocation should dynamically and collaboratively encompass bandwidth, time, beams, and power across multiple satellites.
IV-F Positioning Privacy and Communication Security
[…]
V Conclusion
This paper provides a comprehensive overview of IPAC systems on LEO satellites. By examining the unique characteristics of LEO satellite systems, we demonstrate that the synergistic integration of positioning and communication functionalities can lead to significant mutual enhancement. Despite these benefits, the development of IPAC systems on LEO networks also presents considerable challenges that require further efforts. This remains an open area for research, inviting more in-depth exploration into the promising direction of IPAC systems.
VI Acknowledgments
Fig. 1 was created by Ana Runte, Scientific Illustrator at KAUST.